OpenStudy (anonymous):

A soccer ball kicked on a level field has an initial vertical velocity component of 15.0 meters per second. Assuming the ball lands at the same height from which it was kicked, what is the total time the ball is in the air? [Neglect friction.]

OpenStudy (lasttccasey):

I'm assuming the ball was kicked straight up in the air? If my memory servers me right, dividing by 9.8m/s^2 (gravity) should give you the time it was rising, double that to get falling time, so something in the nature of 3.1 seconds I would say. It has been 3 years since I took physics in high school so hopefully someone more reliable can answer this for sure.

OpenStudy (anonymous):

Since the ball was kicked in air Vertically Therefore: the gravitational Acceleration=-9.8 m/s^2 & V=0 m/s Where U=15.0 m/s Since: V=U+GT therefore:0=15.0-9.8T ,So, T=15/9.8=1.53 Sec Since:the ball took the same distance up & down So the time is = Too therefore:Total (T)=1.53X2=3.06 Sec Appro.

OpenStudy (lasttccasey):

Thank you for the better explination, I rounded to 3.1s but I also got 3.06s to be exact.

OpenStudy (anonymous):

JopHP, what does U and GT stand for?

OpenStudy (anonymous):

U: is the initial Velocity while GT: is the gravitational Acceleration X Time law of motion v=u+at

OpenStudy (anonymous):

\[t=(2u \sin \theta)/(g)\] t=(2*15)/(9.8) because the vertical velocity component is u sin theta. t=3.061 seconds